OpenStudy (anonymous):

PLZ PLZ HELP ME!!!!!!!!!! V-5/3-V is greater than or equal to -3(v-1)

OpenStudy (anonymous):

whr r u stuck

OpenStudy (anonymous):

i juss wunt somebody to work this out for me and i'll be able to see where i went wrong

OpenStudy (anonymous):

um, show me what u did

OpenStudy (anonymous):

\[\frac{ v-5 }{ 3-v } = -3(v-1) \]

OpenStudy (anonymous):

that's it right

OpenStudy (anonymous):

except its greater than or equal

OpenStudy (anonymous):

OK